a ball is projectd from the ground with the speed of 70m/s at theangle 45*with the vertical so that it strikes a verticall ball at the 490/3 m from the point of projection.if the ball return to the point of projection without any colisson with the ground ,then finf coefficent of restution between baal and the wall

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jauneet

Total time of flight

t =2usin45/g =2*70/10√2

                   =7√2

in this time it has to cover 490/3 distance with velocity 70cos45   and  490/3 distance with velocity e70cos45/

7√2 = 490/3*70cos45       + 490/3*e70cos45

7 = 7/3  + 7/3e

1= 1/3 + 1/3e

e=1/2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin